OpenStudy (anonymous):

Need help, very confused? A city government has approved the construction of an $800 million sports arena. Up to $480 million will be raised by selling bonds that pay simple interest at a rate of 7% annually. The remaining amount (up to $640 million) will be obtained by borrowing money from an insurance company at a simple interest rate of 2%. Determine whether the arena can be financed so that the annual interest is $30 million. (If there is no solution, enter NO SOLUTION.) The arena should be financed by selling _____ amount of bonds and borrowing ____ million

OpenStudy (anonymous):

Million in bonds, not amount of bonds, sorry!

OpenStudy (anonymous):

If you can show me step by step that would be great

OpenStudy (amistre64):

can we fund this thing by paying no more than 30mil in interest a year is what is asking for right?

OpenStudy (amistre64):

or maybe its just a simple system of equations x + y = 800 .02x + .07y = 30

OpenStudy (amistre64):

-.02x -.02y = -.02(800) .02x+ .07y = 30 --------------------- .05y = 30 - 16 y = 14/.05 = 1400/5 = 280mil x = 520 mil